OpenStudy (roach):

Joe is paid a weekly salary of $520. He is paid an additional $21 for every hour of overtime he works. This week his total pay was $604. How many hours of overtime did he work? solve it

OpenStudy (anonymous):

520 + 21h = 604 21h = 604 - 520 21h = 84 h = 84/21 h = 4 Joe worked 4 hours overtime.

OpenStudy (roach):

thank you!

OpenStudy (anonymous):

anytime :)